# Learning Patterns from Biological Networks: A Compounded Burr Probability Model

### v2

@@ V.3 Analysis of results
The proposed CBurr and other distributions were implemented using R Software version 4.2.2 on a Windows desktop, equipped with an Intel Core i7-8700 CPU at 3.20 GHz, 6 GB of RAM, and a 64-bit operating system. These models can be easily run and simulated on a personal computer across various platforms. In Table 1, we first provide the statistical properties of the different networks, including their node count, edge count, and descriptive statistics. In particular, the standard deviation (s), mean ($\mu$), and coefficient of variation (s/$\mu$). It is interesting to note that the coefficient of variation for all these datasets is higher than 1, indicating datasets have high variability. Table 1 also presents evidence of the proposed fitting using the Compounded Burr distribution for the node degree distribution in various networks. The parameters $(\gamma,\alpha,c,\lambda)$ of the Compounded Burr distribution were estimated numerically using the “optim” function with the L-BFGS-B algorithm in R software, initialized at $(\gamma,\alpha,c,\lambda)=(1,1,0.5,1.5)$. The advantage of using the “optim” function with the L-BFGS-B algorithm is its ability to efficiently find the optimal parameter values by iteratively optimizing the objective function based on the provided initial values and constraints. The estimated parameter values in Table 1 satisfy the conditions $(\gamma>0,\;\alpha>0,\;c>0,\;\text{and}\;\lambda>-2)$, which are necessary for the complete characterization of the CBurr distribution. Notably, the estimated values of $\gamma$ tend to be higher than the other parameters $\alpha,c,\;\text{and}\;\lambda$, suggesting the scale parameter’s dominant influence in the distribution. Furthermore, the estimated value of the parameter $\lambda$ ranges between $(-1,\infty)$, indicating a wide range of variability in the shape parameter’s values, as observed from Table 1.
@@
In the $\chi^{2}$ test for goodness of fit, the Null Hypothesis $(H_{o})$ is that the observed data is consistent with the predicted values of the fitted distribution. If the p-value of the test is lower than 0.05, then the test rejects the Null Hypothesis; otherwise, we fail to reject $(H_{o})$ Efron and Tibshirani (1994); Davison and Hinkley (1997). When the Bootstrap Chi-square test p-value is higher than $0.9$, it suggests the test provides insufficient evidence against the observed data fits well with the proposed CBurr distribution model. This is further supported by the observation that eight out of twelve datasets in Table 1 have p-values higher than 0.9. The high p-values indicate that there is not enough evidence to reject the null hypothesis, indicating that the observed data is consistent with the proposed CBurr distribution model. Therefore, when the Bootstrap Chi-square value $(p)$ exceeds $0.9$, it provides strong evidence supporting the validity of the proposed CBurr distribution model for the tested biological network data. Table 2 presents statistical measures (RMSE, KLD, and MAE) used to evaluate the performance of the CBurr distribution. The results show that the CBurr distribution generally achieves smaller RMSE and MAE values compared to other distributions, with a few exceptions favoring the power-law cutoff or Burr distribution. Out of twelve datasets, ten of them had the best RMSE performances when using the proposed CBurr distribution model in comparison with others. The Poisson distribution consistently exhibits the poorest performance across all networks. Table 2 displays the highest values of RMSE, KLD, and MAE, which have been highlighted in bold, across all the datasets among the considered distributions. Regarding KLD values, the CBurr distribution consistently yields smaller values compared to other distributions, except in cases where the power-law cutoff or Burr distribution performs better. This indicates a favorable match between the observed degree distribution and the CBurr distribution in most networks. Notably, the Poisson distribution consistently performs worse than other distributions in terms of KLD, RMSE, and MAE. In summary, considering the RMSE, KLD, and MAE values, the CBurr distribution outperforms other distributions across all networks. These results strongly support the plausibility of modeling the observed distribution with the CBurr distribution.
@@
The efficacy of the proposed CBurr distribution can be confirmed by examining the log-log plots of fitted results obtained from various distribution functions. For each network, these plots display the actual frequency distribution, estimated frequency using the CBurr distribution, and estimated frequencies using other distributions such as Pareto Type I, Power Law, Log Normal, Power Law Cutoff, Lomax, Burr, Exponentiated Burr, and MO Burr. The PDF and CDF, along with supports and parameters, are given in Table 3. To make the comparison fair, we implemented them in a similar manner as done in CBurr. The detailed implementation of all the competing methods in the CBurr model is given in https://github.com/ctanujit/CBurr for reproducibility of the results presented in this study. Fig. 5 showcases these log-log graphs for several biological networks considered in this study. Notably, the CBurr distribution consistently yields a better fit as its curve consistently intersects the midpoint of the observed distribution scatter plot, indicating superior performance compared to the other considered distributions. However, in select cases, the power-law cutoff and log-normal distributions exhibit even better curve fitting. As observed in all the plots presented in Fig. 5, the Power Law distributions consistently demonstrate poor performance across the majority of cases. The proposed CBurr distribution offers an effective and efficient approach for modeling the entire degree distribution of real-world networks, including lower-degree nodes, thereby enhancing our understanding of their structural properties.
@@
Table 3: PDF and CDF of competing distributions and their parameters
Name pmf/pdf cdf support
Power-law $Cy^{-\alpha}$ $\sum_{y=1}^{k}Cy^{-\alpha}$ $y=1,2,3,\ldots$
Pareto $\alpha\frac{x_{m}^{\alpha}}{y^{\alpha+1}}$ $\sum_{y=1}^{k}\alpha\frac{x_{m}^{\alpha}}{y^{\alpha+1}}$ $y=1,2,3,\ldots$
@@
Log-normal $\frac{1}{y\sqrt{2\pi\sigma^{2}}}\exp\left[-\frac{(\ln y-\mu)^{2}}{2\sigma^{2}}\right]$ $\frac{1}{2}+\frac{1}{2}\operatorname{erfc}\left(\frac{(\ln y-\mu)}{\sqrt{2}\sigma}\right)$ $y\in(0,\infty)$
Poisson $\frac{e^{-\mu}\mu^{y}}{y!}$ $\sum_{y=0}^{k}\frac{e^{-\mu}\mu^{y}}{y!}$ $y=0,1,2,\ldots$
Power-law Cutoff $Cy^{-\alpha}e^{-\lambda y}$ $\sum_{y=1}^{k}Cy^{-\alpha}e^{-\lambda y}$ $y=1,2,3,\ldots$
@@
Lomax $\frac{\alpha}{\gamma}\left(1+\frac{y}{\gamma}\right)^{-\alpha-1}$ $1-\left(1+\frac{y}{\gamma}\right)^{-\alpha}$ $y>0,\alpha,\gamma>0$
@@
Burr $c\alpha\gamma^{-c}y^{c-1}\left[1+\left(\frac{y}{\gamma}\right)^{c}\right]^{-\alpha-1}$ $1-\left[1+\left(\frac{y}{\gamma}\right)^{c}\right]^{-\alpha}$ $y>0,\alpha,\gamma,c>0$
@@
Exponentiated Burr $\alpha\beta\theta y^{\alpha-1}\frac{\left[1-\left(1+y^{\alpha}\right)^{-\theta}\right]^{\beta-1}}{\left(1+y^{\alpha}\right)^{(\theta+1)}}$ $\left[1-\left(1+y^{\alpha}\right)^{-\theta}\right]^{\beta}$ $y>0,\alpha,\beta,\theta>0$
Burr MO $\frac{\alpha cky^{c-1}\left(1+y^{c}\right)^{-(k+1)}}{\left[1-(1-\alpha)\left(1+y^{c}\right)^{-k}\right]^{2}}$ $\frac{1-\left(1+y^{c}\right)^{-k}}{1-(1-\alpha)\left(1+y^{c}\right)^{-k}}$ $y>0,\alpha,c,k>0$.
@@ VI Discussions
In this article, we have introduced the compounded Burr (CBurr) distribution for modeling the node degree distribution of biological networks. The CBurr distribution is derived from a family of continuous probability distributions, allowing for the inclusion of an additional parameter to enhance the flexibility and efficiency of the Burr XII distribution. Our data analysis and experiments have provided strong evidence supporting the superiority of the proposed CBurr distribution over traditional power law distributions. The CBurr distribution demonstrates better modeling capabilities, capturing the intricate connectivity patterns and characteristics observed in real-world biological networks. By incorporating the CBurr distribution, we are able to represent the entire node degree distribution accurately, considering both highly connected nodes (hubs) and nodes with lower degrees. This comprehensive modeling approach enables a more realistic representation of the network structure and facilitates a deeper understanding of biological systems. In future research, we will explore the applicability of the CBurr distribution in a broader range of network types and investigate its performance across various statistical properties.
In conclusion, the proposed CBurr distribution offers a valuable tool for the effective and efficient modeling of node degree distributions in biological networks. Its flexibility, accuracy, and ability to capture complex connectivity patterns make it a promising choice for analyzing and interpreting biological network data. By conducting extensive analyses of different network structures and characteristics, we aim to gain a deeper understanding of the versatility and robustness of the proposed distribution. We are confident that the CBurr distribution will continue to demonstrate its effectiveness and contribute to the modeling and analysis of diverse real-world networks.
@@
Datasets are available on the network repository https://networkrepository.com/bio.php for public use and are cited in this manuscript. Codes and implementation details are available on https://github.com/ctanujit/CBurr.

### v3

@@ V.3 Analysis of results
The proposed CBurr and other distributions were implemented using R Software version 4.2.2 on a Windows desktop, equipped with an Intel Core i7-8700 CPU at 3.20 GHz, 6 GB of RAM, and a 64-bit operating system. These models can be easily run and simulated on a personal computer across various platforms. In Table 1, we first provide the statistical properties of the different networks, including their node count, edge count, and descriptive statistics. In particular, the standard deviation (s), mean ($\mu$), and coefficient of variation (s/$\mu$). It is interesting to note that the coefficient of variation for all these datasets is higher than 1, indicating that the datasets have high variability (highly skewed). Table 1 also presents evidence of the proposed fitting using the Compounded Burr distribution for the node degree distribution in various networks. The parameters $(\gamma,\alpha,c,\lambda)$ of the Compounded Burr distribution were estimated numerically using the “optim” function with the L-BFGS-B algorithm in R software, initialized at $(\gamma,\alpha,c,\lambda)=(1,1,0.5,1.5)$. The advantage of using the “optim” function with the L-BFGS-B algorithm is its ability to efficiently find the optimal parameter values by iteratively optimizing the objective function based on the provided initial values and constraints. The estimated parameter values in Table 1 satisfy the conditions $(\gamma>0,\;\alpha>0,\;c>0,\;\text{and}\;\lambda>-2)$, which are necessary for the complete characterization of the CBurr distribution.
In our empirical analyses, we may observe the Poisson shift parameter $\lambda$ attaining negative values due to poor parameter initialization, model overflexibility, or numerical instability in the optimization process, especially when the data’s tail behavior is already well captured by the baseline Burr distribution. From a statistical perspective, a negative $\lambda$ reverses the exponential skewing introduced via the Poisson-shifted minimum construction, effectively dampening the influence of the Poisson mechanism. This suggests that the data may not support—or even resist—additional compounding from the skewing layer, particularly when the baseline model (such as the Burr) already accommodates heavy-tailed characteristics observed in biological networks. Rather than signaling a misfit, a negative $\lambda$ indicates an adaptive response of the model, aligning with the underlying data by softening or inverting the shift intensity.
To avoid negative values, one may enforce positivity constraints on $\lambda$ during parameter estimation. Yet, this can lead to under-identified models, where different parameter configurations yield similar likelihoods, making inference unstable. Constrained optimization provides a potential remedy, but at the cost of increased computational burden and sensitivity to starting values or tolerance thresholds. This can be considered as a future scope for further investigation. Notably, the estimated values of $\gamma$ tend to be higher than the other parameters $\alpha,c,\;\text{and}\;\lambda$, signifies that $\gamma$ compresses the distribution, shifting the mass towards smaller values of the degree of a node in the network (for a protein (node), it is the number of direct interaction partners (edges) it has). It suggests that a heavier penalty is imposed for higher degrees. For example, in a PPI network, a high $\gamma$ could suggest that proteins tend to have lower maximum interaction capacity, or that the interaction-generating mechanisms are more constrained in scale. Furthermore, the estimated value of the parameter $\lambda$ ranges between $(-2,\infty)$, indicating a wide range of variability in the shape parameter’s values, as observed from Table 1.
@@
Figure 5: Insights into network connectivity patterns and distribution characteristics are provided by the log-log scale degree distribution of the (top-left) bio-Mouse-Gene, (top-right) Bio HS-LC, (middle-left) Bio-HS-CX, (middle-right) Bio-Grid-Yeast, (bottom-left) Bio-Grid-Warm, and (bottom-right) Bio-Grid-Human networks.
In the $\chi^{2}$ test for goodness of fit, the Null Hypothesis $(H_{o})$ is that the observed data is consistent with the predicted values of the fitted distribution. If the p-value of the test is lower than 0.05, then the test rejects the Null Hypothesis; otherwise, we fail to reject $(H_{o})$ Efron and Tibshirani (1994). When the Bootstrap Chi-square test p-value is higher than $0.9$, it suggests the test provides insufficient evidence against the observed data fits well with the proposed CBurr distribution model. This is further supported by the observation that eight out of twelve datasets in Table 1 have p-values higher than 0.9. The high p-values indicate that there is not a significant deviation between the observed and model-generated degree distributions, confirming the CBurr’s strong goodness-of-fit. The proposed CBurr model captures both the heavy tail and non-monotonic body structure often seen in many of the tested biological network data. A $p$-value less than $0.9$, while not necessarily indicating a poor fit, suggests that the CBurr may not fully capture all aspects of the observed degree distribution, particularly in the body or tail of the distribution. This could be due to model misspecification, parameter instability, or data heterogeneity, such as the presence of multiple network regimes or domain-specific noise. Nonetheless, if the $p$-value remains above conventional thresholds (e.g., $>0.05$), the model is still statistically acceptable, albeit with reduced confidence in its robustness.
@@
Table 2 presents statistical measures (RMSE, KLD, and MAE) used to evaluate the performance of the CBurr distribution. The results show that the CBurr distribution generally achieves smaller RMSE and MAE values compared to other distributions, with a few exceptions favoring the power-law cutoff or Burr distribution. Out of twelve datasets, ten of them had the best RMSE performances when using the proposed CBurr distribution model in comparison with others. The Poisson distribution consistently exhibits the poorest performance across all networks. Table 2 displays the highest values of RMSE, KLD, and MAE, which have been highlighted in bold, across all the datasets among the considered distributions. Regarding KLD values, the CBurr distribution consistently yields smaller values compared to other distributions, except in cases where the power-law cutoff or Burr distribution performs better. This indicates a favorable match between the observed degree distribution and the CBurr distribution in most networks. As depicted in Table 2, CBurr model often behaves similarly to the Burr and cut-off power-law distributions when the Poisson shift parameter $\lambda$ is small, causing the exponential term to approximate $1$ and effectively reducing the model to the standard Burr. Additionally, its survival function structurally resembles the cut-off power-law under certain parameter settings, especially in the tails. Since many biological networks exhibit heavy-tailed but smoothly truncated degree distributions, these models tend to offer similarly good fits. As a result, the added flexibility of the CBurr may not noticeably improve performance unless the data shows more complex, non-monotonic patterns. In summary, considering the RMSE, KLD, and MAE values, the CBurr distribution outperforms other distributions across all networks. These results strongly support the plausibility of modeling the observed distribution with the CBurr distribution.
@@
The efficacy of the proposed CBurr distribution can be confirmed by examining the log-log plots of fitted results obtained from various distribution functions. For each network, these plots display the actual frequency distribution, estimated frequency using the CBurr distribution, and estimated frequencies using other distributions such as power-law, power-law cutoff, and Burr distributions. The PDF and CDF, along with supports and parameters, are given in Table 3. To make the comparison fair, we implemented them in a similar manner to that done in CBurr. Figs. 4 and 5 showcase these log-log graphs for several biological networks considered in this study. Notably, the CBurr distribution consistently yields a better fit as its curve consistently intersects the midpoint of the observed distribution scatter plot, indicating superior performance compared to the other considered distributions. The proposed CBurr distribution offers an effective and efficient approach for modeling the entire degree distribution of real-world networks, including lower-degree nodes, thereby enhancing our understanding of their structural properties.
@@
Table 3: Parameters, PDF, CDF, and comparison of competing distributions with the proposed CBurr distribution
Name pmf/pdf cdf support No. of Remarks
Parameters
Power-law $Cy^{-\alpha}$ $\sum_{y=1}^{k}Cy^{-\alpha}$ $y=1,2,3,\ldots$ 2 Very simple
but inflexible
Pareto $\alpha\frac{x_{m}^{\alpha}}{y^{\alpha+1}}$ $\sum_{y=1}^{k}\alpha\frac{x_{m}^{\alpha}}{y^{\alpha+1}}$ $y=1,2,3,\ldots$ 2 Heavy-tailed
@@
Log-normal $\frac{1}{y\sqrt{2\pi\sigma^{2}}}\exp\left[-\frac{(\ln y-\mu)^{2}}{2\sigma^{2}}\right]$ $\frac{1}{2}+\frac{1}{2}\operatorname{erfc}\left(\frac{(\ln y-\mu)}{\sqrt{2}\sigma}\right)$ $y\in(0,\infty)$ 2 Mean and standard
deviation of log-degrees
Poisson $\frac{e^{-\mu}\mu^{y}}{y!}$ $\sum_{y=0}^{k}\frac{e^{-\mu}\mu^{y}}{y!}$ $y=0,1,2,\ldots$ 1 Light-tailed
Power-law Cutoff $Cy^{-\alpha}e^{-\lambda y}$ $\sum_{y=1}^{k}Cy^{-\alpha}e^{-\lambda y}$ $y=1,2,3,\ldots$ 3 Adds exponential cutoff
to Power-law
@@
Lomax $\frac{\alpha}{\gamma}\left(1+\frac{y}{\gamma}\right)^{-\alpha-1}$ $1-\left(1+\frac{y}{\gamma}\right)^{-\alpha}$ $y>0,\alpha,\gamma>0$ 2 Shifted Pareto
@@
Burr $c\alpha\gamma^{-c}y^{c-1}\left[1+\left(\frac{y}{\gamma}\right)^{c}\right]^{-\alpha-1}$ $1-\left[1+\left(\frac{y}{\gamma}\right)^{c}\right]^{-\alpha}$ $y>0,\alpha,\gamma,c>0$ 3 Captures wide range
of behaviors
@@
Exponentiated Burr $\alpha\beta\theta y^{\alpha-1}\frac{\left[1-\left(1+y^{\alpha}\right)^{-\theta}\right]^{\beta-1}}{\left(1+y^{\alpha}\right)^{(\theta+1)}}$ $\left[1-\left(1+y^{\alpha}\right)^{-\theta}\right]^{\beta}$ $y>0,\alpha,\beta,\theta>0$ 3 Extends Burr with
resilience parameter
Burr MO $\frac{\alpha cky^{c-1}\left(1+y^{c}\right)^{-(k+1)}}{\left[1-(1-\alpha)\left(1+y^{c}\right)^{-k}\right]^{2}}$ $\frac{1-\left(1+y^{c}\right)^{-k}}{1-(1-\alpha)\left(1+y^{c}\right)^{-k}}$ $y>0,\alpha,c,k>0$ 3 Extends Burr with
tilt parameter
Proposed CBurr R.H.S. of Eqn. 16 $1-$ R.H.S. of Eqn. 15 $y>0,\lambda,\alpha,\beta,\theta>0$ 4 Extends Burr with $\lambda$
(Poisson rate), allows skewing,
controls body-tail balance
@@ VI Discussions
This article introduced the CBurr distribution for modeling the node degree distribution of biological networks. The CBurr distribution is derived from a family of continuous probability distributions, allowing for the inclusion of a Poisson-shift parameter to enhance the flexibility and efficiency of the Burr XII distribution. Our data analysis and experiments have provided strong evidence supporting the superiority of the proposed CBurr distribution over traditional power-law distributions. The CBurr distribution demonstrates better modeling capabilities, capturing the intricate connectivity patterns and characteristics observed in real-world heavy-tailed biological networks. By incorporating the CBurr distribution, we are able to represent the entire node degree distribution accurately, considering both highly connected nodes (hubs) and nodes with lower degrees. This comprehensive modeling approach enables a more realistic representation of the network structure and facilitates a deeper understanding of biological systems.
In conclusion, the proposed CBurr distribution offers a valuable tool for the effective and efficient modeling of node degree distributions in biological networks. Its flexibility, accuracy, and ability to capture complex connectivity patterns make it a promising choice for analyzing and interpreting biological network data. By conducting extensive analyses of different network structures and characteristics, we aim to gain a deeper understanding of the versatility and robustness of the proposed distribution. While the CBurr distribution offers high flexibility, it comes with increased computational cost due to the normalization constant and four-parameter estimation. It risks overfitting, especially on smaller or sparse datasets, and its parameters can be difficult to interpret biologically. While excellent for empirical fit, it lacks a generative mechanism rooted in biological network theory. This can be considered a possible future research avenue for understanding the generative mechanism of biological networks using the proposed CBurr model. As future research, exploring CBurr distribution for analyzing social networks, collaboration networks, temporal networks, and web networks will demonstrate its effectiveness and contribute to the modeling and analysis of diverse real-world networks in other areas.
@@
Datasets are collected from the network repository https://networkrepository.com/bio.php. Codes and implementation of the models are available at https://github.com/ctanujit/CBurr.
Acknowledgements.
The authors would like to thank the editor and the two learned reviewers for their constructive discussions and valuable comments.